Nova Company’s total overhead cost at various levels of activity are presented below: Month Machine- Hours Total Overhead Cost

April 46,000 $ 180,540 May 36,000 $ 156,640 June 56,000 $ 204,440 July 66,000 $ 228,340

Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 36,000 machine-hour level of activity is:

Utilities (variable) $ 43,200 Supervisory salaries (fixed) 58,000 Maintenance (mixed) 55,440 Total overhead cost $ 156,640 Nova Company’s management wants to break down the maintenance cost into its variable and fixed cost elements. Required: 1. Estimate how much of the $228,340 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $228,340 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.) 2. Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX. 3. Express the company’s total overhead cost in the form Y = a + bX. 4. What total overhead cost would you expect to be incurred at an activity level of 41,000 machine-hours?

A. Estimate how much of the $228,340 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $228,340 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.) (Do not round intermediate calculations.)

B. Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX. (Do not round your intermediate calculations. Round the "Variable cost per unit" to 2 decimal places.)

C.

Express the company’s total overhead cost in the form Y = a + bX. (Do not round your intermediate calculations. Round the "Variable cost per unit" to 2 decimal places.)

Variable Cost per
Machine-Hour Fixed Cost
Utilities cost
Supervisory salaries cost
Maintenance cost
Total overhead cost

Y= _______ + ________ X

D. What total overhead cost would you expect to be incurred at an activity level of 41,000 machine-hours?
Answer #1

1.Utilities cost is variable, Hence, utilities cost for July = 43,200*66,000/36,000 = $79,200

Supervisory cost is fixed at $58,000

Overhead cost of July = $228,340

Less: Utilities Cost = $79,200

Less: Supervisory Cost = $58,000

Maintenance Cost = $91,140

2.Variable part in maintenance cost using high low method = Difference in cost/Difference in hours

= (91,140-55,440)/(66,000-36,000)

= $1.19 per hour

Fixed Part = 91,140– 1.19*66,000

= $12,600

Hence, equation = 12,600 + 1.19*Number of hours

3.Total overhead cost equation = (12600+58000) + (1.2+1.19)*Number of hours

= 70600 + 2.39x

4.Total overhead cost = 70,600 + 2.39*41,000

= $168,590